This visualization is for questions 2a and 2b. The following dot plots display the distributions of daily high temperatures for the month of February 2006 in Lincoln, Nebraska; San Luis Obispo, Califormia; and Sedona Arizona. 6 70 10 20 40 10 30 40 0 60 70 0 10 2 450 7080 2a. (3 points) Which city tended to have the highest daily high temperatures in that month? Which city tended to have the lowest daily high temperatures in that month? In 1-2 sentences, explain your reasoning 2b. (3 points) Which city had the most day-to-day consistency in daily highest temperatures in that month? Which city tended to have the lowest day-to-day consistency in daily highest temperatures in that month? In 1-2 sentences, explain your reasoning.
Answer #1

(2a) From the dot plots, we can see that San Luis Obispo had the highest daily high temperature of approximately 88 ° F and Lincoln had the lowest daily high temperature of approximately 12 ° F

(2b) The range of daily high temperatures for the three cities are Lincoln = 74 - 12 = 62 ° F, San Luis Obispo = 88 - 52 = 36 ° F and Sedona = 68 - 48 = 20 ° F. Thus, Sedona had the highest day-to-day consistency in the highest daily temperatures and Lincoln had the lowest day-to-day consistency in the highest daily temperatures.